OpenStudy (anonymous):

Can someone check and see if i did this right because i kinda dont understand. Find the solutions to 0 > –(x – 2)3 by graphing y > –(x – 2)3 by hand.

OpenStudy (anonymous):

Solutions (0,8) , (2,0)

OpenStudy (anonymous):

@sourwing

OpenStudy (anonymous):

@love_jessika15

OpenStudy (doc.brown):

If x is zero -(0-2)3 -(-2)3 (+2)(3) =6

OpenStudy (anonymous):

I Dont Understand

OpenStudy (anonymous):

Please Help

OpenStudy (anonymous):

Aww I See it now

OpenStudy (anonymous):

So 6 is the solution

OpenStudy (doc.brown):

Does that graph come from the same question?

OpenStudy (anonymous):

No I Just Graphed It

OpenStudy (doc.brown):

Why do you put a capital letter on every word?

OpenStudy (anonymous):

Thats How I Text Its A Habit !

OpenStudy (doc.brown):

Is it \(-(x-2)^3\)?

OpenStudy (doc.brown):

It looks like everything you write is a "Title of a Movie or Book"

OpenStudy (anonymous):

LOL So for this one Find the solutions to 0 ≥ x^4 – 1 by graphing y ≥ x4 – 1 by hand its ganna be (0^4)-1

OpenStudy (doc.brown):

if x=0 y=-(0-2)^3 -(-2)^3 (+2)^3 =8 Sorry, yes. Your answers are correct.

OpenStudy (anonymous):

So Wat I Did The First Time Was Correct

OpenStudy (doc.brown):

Yes if x=0 then y=0^2-1 = -1

OpenStudy (doc.brown):

It was correct, i thought it was -(x-2) times 3, my mistake

OpenStudy (anonymous):

So The Graphs Are Right To

OpenStudy (doc.brown):

yup

OpenStudy (anonymous):

Oh Okay ! Thanks , Lol

OpenStudy (anonymous):

So can u check this and tell me if its right, feel like im doin it wrong